# Orlicz-Sobolev Transport for Unbalanced Measures on a Graph

### v1

@@ 4.2 Relations to Other Transport Distances
The proof is placed in Appendix §A.8.
Connection with unbalanced Sobolev transport.
Proposition 4.4.
For $N$-function $\Phi(t)=\frac{(p-1)^{p-1}}{p^{p}}t^{p}$ with $1<p<\infty$, and denote $\mathcal{US}_{p,\alpha}$ for the unbalanced Sobolev transport, [^8] See §B.6 for a review on $p$-order UST $\mathcal{US}_{p,\alpha}$. then we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\mathcal{US}_{p,\alpha}(\mu,\nu).$
The proof is placed in Appendix §A.9.
For the limit case [^9] Notice that $\Phi(t)=t$ is not an $N$-function due to its linear growth. It can be considered as the limit $p\to 1^{+}$ of the $N$-function $\Phi(t)=t^{p}$ with $p>1$.: $\Phi(t)=t$
Proposition 4.5 (Limit case for OST).
For $\Phi(t)=t$, and with the same assumptions in Corollary 3.7, then OST yields a closed-form expression as follows:
Moreover, it is equal to the $1$-order unbalanced Sobolev transport (Le et al., 2023).
The proof is placed in Appendix §A.10.
Proposition 4.6 (Limit case for Orlicz-EPT).
For $\Phi(t)=t$, and $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, then Orlicz-EPT is equal to EPT on a graph (Le et al., 2023), i.e.,
$\mathcal{OE}_{\Phi}(\mu,\nu)=\mathrm{KT}(\mu,\nu).$ (22)
The proof is placed in Appendix §A.11.
Proposition 4.7 (Relation of OST and Orlicz-EPT).
Assume that $\Phi(t)=t$, $\omega$ is the length measure on ${\mathbb{G}}$, $\lambda\geq 0$, nonnegative weight functions $w_{1},w_{2}$ are $b$-Lipschitz w.r.t. $d_{{\mathbb{G}}}$, $\alpha=0$, $p=1$. Then we have

### v2

@@ A.2 Detailed Proofs
The proof is completed. ∎
A.2.10 Proof for Proposition 5.4
Proof.
$\hskip-200.0003pt\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+$
$\hskip 80.00012pt\inf_{k>0}\frac{1}{k}\left(1+\int_{{\mathbb{G}}}\frac{(p-1)^{p-1}}{p^{p}}k^{p}b^{p}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right).$ (44)
For convenience, for $k>0$, let
$T(k):=\frac{1}{k}+\frac{(p-1)^{p-1}}{p^{p}}k^{p-1}b^{p}\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x),$
i.e., the objective function of the univariate optimization problem for $\mathcal{OS}_{\Phi,\alpha}$.
We next consider two cases:
$\inf_{k>0}T(k)=\inf_{k>0}\frac{1}{k}=0.$
Consequently, from Equation (A.2.10), we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)$ $=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|=b\,\Big[\int_{{\mathbb{G}}}|\mu(\Lambda(x))-\nu(\Lambda(x))|^{p}\,\omega(\mathrm{d}x)\Big]^{\frac{1}{p}}+\,\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|$
$=\mathcal{US}_{p,\alpha}(\mu,\nu).$
$\lim_{k\to 0^{+}}T(k)=\lim_{k\to+\infty}T(k)=+\infty.$
Therefore, Equation (A.2.10), we have
Additionally, we have
Therefore, by plugging this value of $k_{0}$ into $T$, we have
$\hskip-10.00002ptT(k_{0})$ $=\frac{1}{k_{0}}\left(1+\frac{(p-1)^{p-1}}{p^{p}}k_{0}^{p}b^{p}\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right)$
$\hskip 1.99997pt\left(1+\frac{(p-1)^{p-1}}{p^{p}}\frac{1}{\frac{(p-1)^{p}}{p^{p}}b^{p}\left(\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right)}b^{p}\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right)$
Thus, by plugging this value of $T(k_{0})$ into Equation (45), we obtain
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)$ $=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+b\left(\int_{{\mathbb{G}}}\left|\mu(\Lambda(x))-\nu(\Lambda(x))\right|^{p}\omega(\text{d}x)\right)^{\frac{1}{p}}$
$=\mathcal{US}_{p,\alpha}(\mu,\nu).$
The proof is completed.
A.2.11 Proof for Proposition 5.5
Proof.
Following Corollary 4.3, we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+\inf_{k>0}\frac{1}{k}\left(1+\hskip-3.00003pt\sum_{e\in E}\hskip-1.00006ptw_{e}\Phi\!\left(kb\left|\mu(\gamma_{e})-\nu(\gamma_{e})\right|\right)\right).$
For $\Phi(t)=t$, then we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)$ $=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+\inf_{k>0}\frac{1}{k}\left(1+\hskip-3.00003pt\sum_{e\in E}\hskip-1.00006ptw_{e}kb\left|\mu(\gamma_{e})-\nu(\gamma_{e})\right|\right)$
Hence, the proof is completed. ∎
A.2.12 Proof for Proposition 5.6
Proof.
From Equation (8), we have
For $\Phi(t)=t$, we further have
[…]
Hence, we have
$\mathcal{OE}_{\Phi}(\mu,\nu)$ $=\left(\mu({\mathbb{G}})+\nu({\mathbb{G}})\right)\left(\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu})-b\lambda\right)$
$=\mathrm{KT}(\mu,\nu).$
The proof is completed.
A.2.13 Proof for Proposition 5.7
Proof.
For $\Phi(t)=t$, $p=1$, from Theorem 4.2, we have
$=\mathcal{US}_{p,\alpha}(\mu,\nu).$
Additionally, for $\Phi(t)=t$, from Proposition 5.6, we have
$\mathcal{OE}_{\Phi}(\mu,\nu)=\mathrm{KT}(\mu,\nu).$
With additional assumptions that $\lambda\geq 0$ and the nonnegative weight functions $w_{1},w_{2}$ are $b$-Lipschitz w.r.t. $d_{\mathbb{G}}$, then by applying [40, Corollary 3.2], we have
Consequently, for $\alpha=0$, and the length measure $\omega$ on $\mathbb{G}$}, then following [40, Proposition 5.2], we have